A man invested a sum of money at 3% and three times that amount at 2%. If the total interest received on the investments was Php 270, find the sum invested at each rate
Answer by uma(370)

Let the amount invested at 3% be x
So the amount invested at 2% = 3x
Interest on x at 3% for an year = PTR
= x*0.03*1
= 0.03x
Interest on 3x at 2% for an year = PTR
= 3x * 0.02 * 1
= 0.06x
Total interest = 270
==> 0.03x + 0.06x = 270
==> 0.09x = 270
==> 0.09x/0.09 = 270/0.09
==> x = 3000
So 3x = 9000
He invested 3000 at 3% and 9000 at 2%

Good Luck!!!